OpenStudy (anonymous):

What is the approximate area of a circle with a radius of 28 centimeters? Use π ≈ 3.14. A. 175.8 square centimeters B. 4923.5 square centimeters C. 87.9 square centimeters D. 2461.8 square centimeters

OpenStudy (anonymous):

3.14 x pi^2

OpenStudy (anonymous):

Area of a circle is: \(\pi r^2\) So what's 28 * 28?

OpenStudy (anonymous):

So can you figure out 3.14 x 28^2? @~*Ashtyn*~

OpenStudy (anonymous):

I'm confused -_-

OpenStudy (anonymous):

First, what is 28^2?

OpenStudy (anonymous):

56

OpenStudy (anonymous):

Ummmm... @AnswerMyQuestions

OpenStudy (anonymous):

the square of a number is the number times itself 28 times 28

OpenStudy (anonymous):

go back to formula to find area of a circle pi (3.14) times the radius (28^2) to solve a= 3.14 times 28 times 28

OpenStudy (anonymous):

So 28*28 is what you want me to figure out ?

OpenStudy (anonymous):

that would be the radius squared that is needed in the formula

OpenStudy (anonymous):

28*28=784

OpenStudy (anonymous):

yes now put that number in the formula times pi (3.14)

OpenStudy (anonymous):

formula again to find area of circle is pi times radius squared